OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 7 and a9 = 55 ?

OpenStudy (anonymous):

55-7 48 ---- = ------ =24 9-7 2 d=24

OpenStudy (anonymous):

an=a1+(n-1)d

OpenStudy (anonymous):

I'm pretty sure its not 24